Tiling a rectangle kitchen back-splash measures 60 inches by 12 inches needs to be tiled, each box contains 100 tiles that measure 1 inch by 11 inch how many boxes of tiles are needed?
Answer by richwmiller(17219)

60*12=720 square inches
if the tiles measure 1 by 1 inch they have an area of 1 sq inch
so 720 tiles are needed
720/100=7.2 so you need to buy 8 boxes
if it is 1 by 11 inchesthen an area of 11 sq inches per tile
720/11=65 5/11
which is less than one box so one box is needed